OpenStudy (anonymous):

calculate the slope of the line between points (7,-6) and (-3,-2) I am sooo lost

OpenStudy (anonymous):

formula of slope of line: \[(y _{1}-y _{2})/(x _{1}-x _{2})\] so, apply the formula here: [-6-(-2)]/[7-(-3)]= -4/10=-2/5

sam (.sam.):

\[\huge Slope=\frac{y_{2}-y_{1}}{x_{2}-x_{1}}\]

OpenStudy (anonymous):

@. how did u do the font? its so good! for the problem, follow the above eqn putting x1,y1,x2,y2 perfectly and get the answer